OpenStudy (juana02):

can someone help me find the solution to 9-(-4)? i need to know the steps on how to solve

OpenStudy (zzr0ck3r):

\(9-(-4) = 9+4\)

OpenStudy (jhannybean):

Think of it like : \(9-(-1)(4)\)

Directrix (directrix):

@juana02 Subtraction is addition of negative numbers to my thinking. Change the sign of the second number and add it to the first. 9 - (-4) = 9 + (+4) = 13.

OpenStudy (jhannybean):

When you multiply two negative values, you get a positive value.

OpenStudy (zzr0ck3r):

I like to think of the two negative signs, - - making a positive sign +

OpenStudy (zzr0ck3r):

just turn one and put it on the other...

OpenStudy (accessdenied):

The above methods work, and probably the best way to think about it for practical use. If it helps you learning this, another way to think of it as well is using a number line. A difference of two numbers ends up being the distance between the two on the number line!